given the following function, find the indicated values
f(x)=-7x+1
a) f(o)
b) f(-1)
c) f(3)

Answer by mathslover(157)


f(x)=-7x+1
to get the respective values of f(x) substitute the x with the value in the
function expression
a) f(0) = -7*0 +1 = 1
b) f(-1) = -7 * (-1) +1 = 8
c) f(3)= -7* 3 + 1 = 20